# Fan distributions via Tverberg partitions and Gale duality

### v2

@@ 5.1. Typical Point Sets
In the real case, let $n=(r-1)(d+m+1)+k+1$, where $10\leq k\leq r-$ 1, and assume that $\ell>2+\frac{k}{r-1}$. Suppose that $a_{1},\ldots,a_{n},a_{n+\ell-1}$ are points in strong general position in $\mathbb{R}^{d+\ell-1}$, and let $x_{1},\ldots,x_{n+\ell-1}$ be their Gale dual in $\mathbb{R}^{n-d-1}$. These points are all in general linear position in $\mathbb{R}^{n-d-1}$ (i.e., any $n-d-1$ of them linearly span $\mathbb{R}^{n-d-1}$), so in particular none of the $x_{i}$ is zero. Now consider the sequence $a_{1}^{\prime}=(a_{,}1),\ldots,a_{n+\ell-1}^{\prime}=(a_{n+\ell-1},1)$ in $\mathbb{R}^{d+\ell}$, and append to this sequence any vector $a_{n+\ell}^{\prime}$ in $\mathbb{R}^{d+\ell}$ such that $a_{1}^{\prime},\ldots,a^{\prime}_{n+\ell}$ affinely span $\mathbb{R}^{d+\ell}$. We now claim that, for a suitable choice of basis, the Gale transform of the $a_{i}^{\prime}$ is $x_{1},\ldots,x_{n+\ell-1},x_{n+\ell}$, where $x_{n+\ell}=0$. To see this, let $A$ be the $(d+\ell)\times(n+\ell-1)$ matrix in the construction of the Gale transform of $a_{1},\ldots,a_{n+\ell-1}$, and let $b_{1},\ldots,b_{n-d-1}\in\mathbb{R}^{n+\ell-1}$ be a choice of basis for $\ker A$ so that $x_{1},\ldots,x_{n+\ell-1}$ are the columns of the matrix $B$ whose rows are the $b_{i}$. Letting $A^{\prime}$ be the matrix in the construction of the Gale transform of the $a_{i}^{\prime}\in\mathbb{R}^{d+\ell}$, we may choose $b_{1}^{\prime}=(b_{1},0),\ldots,b_{n-d-1}^{\prime}=(b_{n-d-1},0)$ as a basis for $\ker A^{\prime}$. Thus the Gale transform of $a_{1}^{\prime},\ldots,a^{\prime}_{n+\ell}$ in $\mathbb{R}^{n+\ell}$ is $x_{1},\ldots,x_{n+\ell-1},x_{n+\ell}=0$ as claimed. As any $n-d-1$ of $x_{1},\ldots,x_{n+\ell-1}$ linearly span $\mathbb{R}^{n-d-1}$ and $x_{1},\ldots,x_{n+\ell-1}$ sums to zero, it follows easily that $x_{1},\ldots,x_{n+\ell}$ affinely span $\mathbb{R}^{n-d-1}$.
@@
We now show that there exists an $m$-coloring of $X=\{x_{1},\ldots,x_{n+\ell}\}$ which cannot be equidistributed by any $r$-fan in $\mathbb{R}^{n-d-1}$ of codimension $r-2$. Namely, let $X=X_{1}\sqcup\cdots\sqcup X_{m}$ be an $m$-coloring such that $|X_{i}|=r-1$ for all $i\geq 2$ and $x_{n+\ell}=0$ is in $X_{m}$. Now suppose that $F_{r}=\cup_{j=1}^{r}B_{j}$ is an $r$-fan of codimension $r-2$ which equidistributes this coloring. As $|\operatorname{\mathrm{Int}}(B_{j})\cap X_{i}|\leq|X_{i}|/r$ for all $i\in[m]$ and $j\in[r]$, all the points of $X\setminus X_{1}$ must lie on the center $C$ of $F_{r}$. In particular, $F_{r}$ is linear. Now let $f\colon\Delta_{n+\ell-1}\rightarrow\mathbb{R}^{d+\ell}$ be the linear map determined by the $a_{i}^{\prime}$ and let $[n+\ell]=C_{1}\sqcup\cdots\sqcup C_{m}$ be the $m$-coloring of the vertex set of $\Delta_{n+\ell-1}$ corresponding to the $m$-coloring of $X$. Applying part (b) of Lemma 3.1, there is a proper Tverberg $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that each $I_{j}:=\operatorname{\mathrm{Vert}}(\sigma_{j})$ is in bijective correspondence with $X\cap\operatorname{\mathrm{Int}}(B_{j})$ and $[n+\ell-1]\setminus\sqcup_{j=1}^{r}I_{j}$ is in bijective correspondence with $C\cap X$. Thus $|C_{i}\cap\operatorname{\mathrm{Vert}}(\sigma)|=|X_{i}\cap\operatorname{\mathrm{Int}}(B_{j})|$ for all $j\in[r]$ and all $i\in[m]$. Thus each $I_{j}$ is contained in the color class $C_{1}\subset[n+\ell-1]$, which consists of $n-(m-1)(r-1)=(r-1)(d+2)+k+1$ points. Letting $\Delta_{C_{1}}$ be the simplex generated by $C_{1}$, we have that the image of the restriction $f\colon\Delta_{C_{1}}\rightarrow\mathbb{R}^{d+\ell}$ of $f$ has image in $\mathbb{R}^{d+\ell-1}\times\{1\}$. Since the points $a_{1}^{\prime}=(a_{1},1),\ldots,a_{n+\ell-1}^{\prime}=(a_{n+\ell-1},1)$ are in strong general position in $\mathbb{R}^{d+\ell-1}\times\{1\}$ and $k<(r-1)(\ell-2)$, we have
$\dim(\cap_{j=1}^{r}\operatorname{\mathrm{Aff}}(f(\sigma_{j})))=\sum_{j=1}^{r}\operatorname{\mathrm{codim}}(\operatorname{\mathrm{Aff}}(f(\sigma_{j}))\geq r(d+\ell)-[(r-1)(d+2)+k+1]>d+\ell-1,$
@@
so $\cap_{j=1}^{r}\operatorname{\mathrm{Aff}}(f(\sigma_{j}))=\emptyset$ and therefore $(\sigma_{1},\ldots,\sigma_{r})$ is not a Tverberg $r$-tuple for $f$. Thus there is no $r$-fan of codimension $r-2$ which equidistributes the $m$-coloring of $X$.
@@
The proof in the complex case when $r=2$ is analogous. Let $n=2(d+1)+m+k$ where $k\in\{0,1\}$. As before, there exist $n+\ell$ complex affinely spanning points $a_{1},\ldots,a_{n+\ell}$ in $\mathbb{C}^{d+\ell}$ whose Gale dual $x_{1},\ldots,x_{n+\ell}$ in $\mathbb{C}^{n-d-1}$ satisfies $x_{n+\ell}=0$ and for which $a_{1},\ldots,a_{n+\ell-1}$ are in strong general position in $\mathbb{C}^{d+\ell-1}\times\{1\}$. In particular, $X$ complex affinely spans $\mathbb{C}^{n-d-1}$. As before, we choose an $m$-coloring of $X=\{x_{1},\ldots,x_{n+\ell}\}$ so that $|X_{i}|=1$ for all $2\leq i\leq m$ and $0\in X_{m}$. If $F_{2}$ were a complex regular $2$-fan which equidistributes the $m$-coloring, then $F_{2}$ would be linear and so by Remark 3.3 the linear map $f\colon\Delta_{n+\ell-1}\rightarrow\mathbb{C}^{d+\ell}$ determined by the $a_{i}$ would have a Tverberg pair $(\sigma_{1},\sigma_{2})$ where each $\sigma_{i}$ does not contain $n+\ell$. As $a_{1},\ldots,a_{n+\ell-1}$ are in strong general position in $\mathbb{C}^{d+\ell-1}\times\{1\}$, as before the condition that $k<2(\ell-2)$ gives $\operatorname{\mathrm{codim}}(\operatorname{\mathrm{Aff}}(f(\sigma_{1})\cap\operatorname{\mathrm{Aff}}(f(\sigma_{2}))>2(d+\ell-1)$. We conclude as before that no equidistributing complex regular $2$-fan exists. ∎
@@
Let $c=(r-1)(m+1)$ and let $d=(r-2)s+t+c$ where $s\geq 1$ and $t\in\{0,\ldots,r-2\}$. Letting $n=d+s+1$ and applying Theorem 1.1, we have that any $m$-coloring of any affinely spanning set of $n=(r-1)(s+m+1)+t+1$ points in $\mathbb{R}^{n-s-1}=\mathbb{R}^{d}$ can be equidistributed by an $r$-fan in $\mathbb{R}^{d}$ of codimension $r-2$. Thus $n(r,m,d)\geq n$. On the other hand, Proposition 5.3 shows that $n(r,m,d)\leq n+\ell-1$ so long as $\ell>2+\frac{t}{r-1}$. But $20\leq t\leq r-$ 2, so $n(r,m,d)\leq n+2$. ∎

### v3

@@ 8.1. Typical Point Sets
In the real case, let n=(r1)(d+m+1)+k+1n=(r-1)(d+m+1)+k+1 where $k\geq 0$ and assume that $\ell>2+\frac{k+2}{r-2}$. Suppose that $a_{1},\ldots,a_{n},a_{n+\ell-1}$ are points in strong general position in $\mathbb{R}^{d+\ell-1}$ and let $x_{1},\ldots,x_{n+\ell-1}$ be their Gale dual in $\mathbb{R}^{n-d-1}$. Since the $x_{i}$ linearly span $\mathbb{R}^{n-d-1}$ and sum to zero they also affinely span $\mathbb{R}^{n-d-1}$. Consider the sequence $a_{1}^{\prime}=(a_{,}1),\ldots,a_{n+\ell-1}^{\prime}=(a_{n+\ell-1},1)$ in $\mathbb{R}^{d+\ell}$ and append to this sequence any point $a_{n+\ell}^{\prime}$ in $\mathbb{R}^{d+\ell}$ such that $a_{1}^{\prime},\ldots,a^{\prime}_{n+\ell}$ affinely spans $\mathbb{R}^{d+\ell}$. We now claim that, for a suitable choice of basis, the Gale transform of the $a_{i}^{\prime}$ is $x_{1},\ldots,x_{n+\ell-1},x_{n+\ell}$, where $x_{n+\ell}=0$. To see this, let $A$ be the $(d+\ell)\times(n+\ell-1)$ matrix in the construction of the Gale transform of $a_{1},\ldots,a_{n+\ell-1}$, and let $b_{1},\ldots,b_{n-d-1}\in\mathbb{R}^{n+\ell-1}$ be a choice of basis for $\ker A$ so that $x_{1},\ldots,x_{n+\ell-1}$ are the columns of the matrix $B$ whose rows are the $b_{i}$. Letting $A^{\prime}$ be the $(d+\ell+1)\times(n+\ell)$ matrix in the construction of the Gale transform of the $a_{i}^{\prime}\in\mathbb{R}^{d+\ell}$, we may choose $b_{1}^{\prime}=(b_{1},0),\ldots,b_{n-d-1}^{\prime}=(b_{n-d-1},0)$ as a basis for $\ker A^{\prime}$. Thus the Gale transform of $a_{1}^{\prime},\ldots,a^{\prime}_{n+\ell}$ in $\mathbb{R}^{n-d-1}$ is $x_{1},\ldots,x_{n+\ell-1},x_{n+\ell}=0$.
@@
We now show that there exists an $m$-coloring of $X=\{x_{1},\ldots,x_{n+\ell}\}$ which cannot be equidistributed by any conical $r$-fan in $\mathbb{R}^{n-d-1}$. Namely, let $X=X_{1}\sqcup\cdots\sqcup X_{m}$ be an $m$-coloring such that $|X_{i}|=r-1$ for all $i\geq 2$ and $x_{n+\ell}=0$ is in $X_{m}$. Now suppose that $F_{r}=\cup_{j=1}^{r}B_{j}$ is a conical $r$-fan which equidistributes this coloring. As $|\operatorname{\mathrm{Int}}(B_{j})\cap X_{i}|\leq|X_{i}|/r$ for all $i\in[m]$ and $j\in[r]$, all the points of $X\setminus X_{1}$ must lie in the center $C$ of $F_{r}$. In particular, $F_{r}$ is linear. Now let $f\colon\Delta_{n+\ell-1}\rightarrow\mathbb{R}^{d+\ell}$ be the linear map determined by the $a_{i}^{\prime}$ and let $n+\ell]=C_{1}\sqcup\cdots\sqcup C_{m}$ be the $m$-coloring of the vertex set of $\Delta_{n+\ell-1}$ corresponding to the $m$-coloring of $X$. By Lemma 3.1 there is a proper Tverberg $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that each $I_{j}:=\operatorname{\mathrm{Vert}}(\sigma_{j})$ corresponds to $X\cap\operatorname{\mathrm{Int}}(B_{j})$ and $n+\ell-1]\setminus\sqcup_{j=1}^{r}I_{j}$ corresponds to $C\cap X$. In particular, $|C_{i}\cap\operatorname{\mathrm{Vert}}(\sigma_{j})|=|X_{i}\cap\operatorname{\mathrm{Int}}(B_{j})|$ for all $j\in[r]$ and all $i\in[m]$. Thus each $I_{j}$ is contained in the color class $C_{1}$, which consists of $n+\ell-(m-1)(r-1)=(r-1)(d+2)+k+\ell+1$ points. Letting $\Delta_{C_{1}}$ be the simplex generated by $C_{1}$, we have that the image of the restriction $f\colon\Delta_{C_{1}}\rightarrow\mathbb{R}^{d+\ell}$ of $f$ is contained in $\mathbb{R}^{d+\ell-1}\times\{1\}$. As the points $a_{1}^{\prime}=(a_{1},1),\ldots,a_{n+\ell-1}^{\prime}=(a_{n+\ell-1},1)$ are in strong general position in $\mathbb{R}^{d+\ell-1}\times\{1\}$ and $k<(r-1)(\ell-2)$-$\ell$, we have
$\operatorname{\mathrm{codim}}(\cap_{j=1}^{r}\operatorname{\mathrm{Aff}}(f(\sigma_{j})))$ $=\sum_{j=1}^{r}\operatorname{\mathrm{codim}}(\operatorname{\mathrm{Aff}}(f(\sigma_{j}))$
$=r(d+\ell)-\sum_{j=1}^{r}|I_{j}|\geq r(d+\ell)-[(r-1)(d+2)+k+\ell+1]>d+\ell-1.$
@@
Thus $\cap_{j=1}^{r}\operatorname{\mathrm{Aff}}(f(\sigma_{j}))=\emptyset$, contradicting that $(\sigma_{1},\ldots,\sigma_{r})$ is a Tverberg $r$-tuple for $f$. Therefore there is no conical $r$-fan in $\mathbb{R}^{n-d-1}$ which equidistributes the $m$-coloring of $X$.
@@
The proof in the complex case when $r=2$ is analogous. Let $n=2(d+1)+m+k$ where $k\geq 0$. As before, there exist $n+\ell$ complex affinely spanning points $a_{1},\ldots,a_{n+\ell}$ in $\mathbb{C}^{d+\ell}$ whose Gale dual $x_{1},\ldots,x_{n+\ell}$ in $\mathbb{C}^{n-d-1}$ satisfies $x_{n+\ell}=0$ and for which $a_{1},\ldots,a_{n+\ell-1}$ are in strong general position in $\mathbb{C}^{d+\ell-1}\times\{1\}$. In particular, $X$ complex affinely spans $\mathbb{C}^{n-d-1}$. As before, we choose an $m$-coloring of $X$ so that $|X_{i}|=1$ for all $2\leq i\leq m$ and $0\in X_{m}$. If $F_{2}$ were a complex regular $2$-fan which equidistributes the $m$-coloring, then $F_{2}$ would be linear and so by Remark 3.3 the linear map $f\colon\Delta_{n+\ell-1}\rightarrow\mathbb{C}^{d+\ell}$ determined by the $a_{i}$ would have a Tverberg tuple $(\sigma_{1},\sigma_{2})$ where each $I_{j}:=\operatorname{\mathrm{Vert}}(\sigma_{j})$ is contained in $C_{1}$. Thus $f|_{\Delta_{C_{1}}}\colon\Delta_{C_{1}}\rightarrow\mathbb{C}^{d+\ell}$ maps to $\mathbb{C}^{d+\ell-1}\times\{1\}$. As $a_{1},\ldots,a_{n+\ell-1}$ are in strong general position in $\mathbb{C}^{d+\ell-1}\times\{1\}$, the condition that $k<\ell-3$ gives $\operatorname{\mathrm{codim}}(\operatorname{\mathrm{Aff}}(f(\sigma_{1}\cap\sigma_{2}))>2(d+\ell-1)$. We conclude as before that no equidistributing complex regular $2$-fan exists. ∎
@@
Let c=(r1)(m+1)c=(r-1)(m+1) and let d=(r2)s+t+cd=(r-2)s+t+c, where $m\geq 2$, $s\geq 1$, and $t\in\{0,\ldots,r-3\}$. Letting $n=d+s+1$ and applying Theorem 1.1, we have that any $m$-coloring of any affinely spanning set of n=(r1)(s+m+1)+t+1n=(r-1)(s+m+1)+t+1 points in $\mathbb{R}^{n-s-1}=\mathbb{R}^{d}$ can be equidistributed by a conical $r$-fan in $\mathbb{R}^{d}$. Thus $n(r,m,d)\geq n$. On the other hand, Proposition 8.4 shows that $n(r,m,d)\leq n+\ell-1$ so long as $\ell>2+\frac{t+2}{r-2}$. But $0\leq t\leq r-3$, so $n(3,m,d)\leq n+4$ and $n(r,m,d)\leq n+3$ when $r>3$. ∎